# Aggregation Models with Optimal Weights for Distributed Gaussian Processes

### v1

@@ 5.1.2 Prediction
Figure 4 compares three metrics for exact GPs: computational time, root mean squared error (RMSE), and negative log predictive density (NLPD) of all aggregation models as a function of the number of training points per GP expert. While grBCM preforms the best on the mean of RMSE and NLPD, it requires more computational time due to the larger size of the augmented local training datasets and has larger standard deviations of the metrics. Our method, GP with optimal weights, provides comparable means for these metrics with smaller standard deviations, indicating greater stability and less sensitivity compared to grBCM. The boxplots of RMSE and NLPD for exact GPs are presented in Figure 9 and Figure 10 in Appendix. These boxplots highlight the stability of our method, especially as the number of experts $M$ increases.
Figure 5 compares three metrics as a function of the number of training points per GP expert for SVGP across all aggregation models. For SVGP, PoE, gPoE, grBCM, significantly outperform the full SVGP, BCM, and rBCM. The boxplots of RMSE and NLPD for SVGP are shown in Figure 11 and Figure 12 in Appendix. These boxplots demonstrate that PoE, gPoE, grBCM, and our method all achieve the smallest means and variabilities for these metrics.
5.2 UCI regression
Figure 6: Comparison of the aggregation models for SVGP on the UCI datasets.
In this section, we evaluate the aggregation models on five real datasets from the UCI repository (Asuncion et al.,, 2007): PoleTele, Elevators, Kin40k, KEGG, and Protein. Each dataset is split into training and test sets in an 80:20 ratio. For the smaller datasets (PoleTele and Elevators, each around $15,000$ samples), the training datasets are divided into $M=4,10,20$ experts. For the larger datasets (Kin40k, KEGG, and Protein, each around $45,000$ samples), the training datasets are divided into $M=8,16,64$ experts. We use SVGP with $m=128$ inducing points and RBF kernel defined in eq. 46. All aggregation models utilize FedAvg with the Adam optimizer, using the same initial hyperparameters over 100 iterations for distributed training. Results are averaged over 10 different seeds. Figure 6 presents the performance metrics of the aggregation models on these five datasets. BCM performs the worst, particularly as the number of experts $M$ increases. While grBCM performs well for RMSE on most datasets, it has significantly worse NLPD on PoleTele. Although our method, distributed SVGP with optimal weights, does not always perform the best, it consistently provides stable and comparable predictions in less time than grBCM. Detailed metric values are reported in Table 2 and Table 3 in Appendix.
6 Conclusion
@@
In this work, we introduced an algorithm to enhance GP regressions with inducing points on sparse grids using the optimized combination technique (OptiCom). Building on this foundation, we proposed a novel aggregated prediction algorithm suitable for both distributed exact GP and SVGP. This method leverages predictions from local models and optimally combines them using weights derived from solving a linear system. Additionally, it inherently incorporates correlations between experts through the computation of optimal weights. The proposed methods have a computational complexity requiring $\mathcal{O}(M^{3})$ additional time compared to PoE family and BCM family for $M$ GP experts. However, this overhead is minimal in practice, as the number of experts $M$ is typically small compared to the size of the local datasets. Through extensive experiments and comparisons with state-of-the-art aggregation models, we demonstrated the competitiveness and stability of our proposed method. Our algorithm consistently provides stable and accurate predictions, offering a robust solution for distributed GP regression tasks.
Acknowledgments and Disclosure of Funding
@@ A.2 Algorithms
[…]
@@
16:  $\mathbf{c}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
@@
3:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{S}:=\text{zeros}(n_{t},n_{t})$
@@
12:  $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*}):=\sigma_{\epsilon}^{-2}\mathbf{m}$
13:  $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*}):=\mathbf{S}$
@@ Appendix B Distributed GP with optimal weights
B.1 Algorithms
@@
3:  $\mathbf{X}_{c}:=\{\mathbf{x}_{i}^{(c)}\}_{i=1}^{M}$, $\mathbf{x}_{i}^{(c)}\in\mathbf{X}_{i}$ is randomly selected
4:  $\mathbf{M}^{(c)}=\big{[}k(\mathbf{Z},\mathbf{Z})+\sigma_{\epsilon}^{-2}k(\mathbf{Z},\mathbf{X}_{c})k(\mathbf{X}_{c},\mathbf{Z})\big{]}$
5:  $\mathbf{A}:=\text{zeros}(M,M)$
@@
18:  $\bm{\beta}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
@@
4:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{s}:=\text{zeros}(n_{t})$
@@
10:  $\mu_{\mathcal{A}}(\mathbf{X}^{*}):=\sigma_{\epsilon}^{-2}\mathbf{m}$
11:  $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*}):=\mathbf{s}$
@@
3:  $\mathbf{X}_{c}:=\{\mathbf{x}_{i}^{(c)}\}_{i=1}^{M}$, $\mathbf{x}_{i}^{(c)}\in\mathbf{X}_{i}$ is randomly selected
4:  $\mathbf{A}:=\text{zeros}(M,M)$
@@
17:  $\bm{\beta}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
@@
4:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{s}:=\text{zeros}(n_{t})$
@@
10:  $\mu_{\mathcal{A}}(\mathbf{X}^{*}):=\mathbf{m}$
11:  $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*}):=\mathbf{s}$
@@ C.1 Boxplots of the hyperparameter estimates
Figure 8: Boxplots of the optimized hyperparemeter estimates on the $n=10^{4}$ dataset.
C.2 Boxplots of the metrics
Figure 9: RMSE of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
Figure 10: NLPD of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
Figure 11: RMSE of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
Figure 12: NLPD of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
C.3 Tables of the UCI datasets
Table 2: Results of the aggregation models for SVGP on the UCI datasets of size $\sim$15K. Best and worst performances are highlighted in yellow and red, respectively.
Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
PoleTele 4 Time 1.588 $\pm$ 0.044 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.910 $\pm$ 0.091 3.939 $\pm$ 0.099
RMSE 16.416 $\pm$ 1.032 41.878 $\pm$ 0.315 41.878 $\pm$ 0.315 44.255 $\pm$ 0.877 41.709 $\pm$ 0.257 16.772 $\pm$ 2.289 35.596 $\pm$ 0.267
NLPD 22.096 $\pm$ 2.964 17.859 $\pm$ 0.245 7.403 $\pm$ 0.061 7.845 $\pm$ 0.174 7.374 $\pm$ 0.050 65.186 $\pm$ 17.279 6.293 $\pm$ 0.052
10 Time 1.588 $\pm$ 0.044 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.826 $\pm$ 0.008 2.121 $\pm$ 0.005
RMSE 48.678 $\pm$ 10.933 41.739 $\pm$ 0.277 41.739 $\pm$ 0.277 44.374 $\pm$ 1.066 41.709 $\pm$ 0.257 17.031 $\pm$ 1.645 30.156 $\pm$ 0.200
NLPD 189.683 $\pm$ 58.036 39.704 $\pm$ 0.529 7.455 $\pm$ 0.054 7.956 $\pm$ 0.210 7.451 $\pm$ 0.050 127.538 $\pm$ 24.461 5.387 $\pm$ 0.039
20 Time 1.588 $\pm$ 0.044 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.944 $\pm$ 0.031 1.104 $\pm$ 0.018
RMSE 48.678 $\pm$ 10.933 41.713 $\pm$ 0.263 41.713 $\pm$ 0.263 42.766 $\pm$ 1.073 41.709 $\pm$ 0.257 17.065 $\pm$ 2.667 25.617 $\pm$ 0.161
NLPD 189.683 $\pm$ 58.036 122.149 $\pm$ 1.430 9.972 $\pm$ 0.079 10.315 $\pm$ 0.347 9.972 $\pm$ 0.077 174.421 $\pm$ 54.154 6.124 $\pm$ 0.048
Elevators 4 Time 1.616 $\pm$ 0.036 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.551 $\pm$ 0.012 4.893 $\pm$ 0.025
RMSE 0.249 $\pm$ 0.004 0.253 $\pm$ 0.005 0.253 $\pm$ 0.005 0.254 $\pm$ 0.005 0.253 $\pm$ 0.005 0.255 $\pm$ 0.005 0.253 $\pm$ 0.005
NLPD 0.846 $\pm$ 0.002 0.792 $\pm$ 0.002 0.822 $\pm$ 0.002 0.822 $\pm$ 0.002 0.822 $\pm$ 0.002 0.815 $\pm$ 0.002 0.822 $\pm$ 0.002
10 Time 1.616 $\pm$ 0.036 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 3.058 $\pm$ 0.008 2.263 $\pm$ 0.008
RMSE 0.255 $\pm$ 0.006 0.253 $\pm$ 0.005 0.253 $\pm$ 0.005 0.324 $\pm$ 0.040 0.253 $\pm$ 0.005 0.259 $\pm$ 0.004 0.253 $\pm$ 0.005
NLPD 0.847 $\pm$ 0.002 0.786 $\pm$ 0.002 0.822 $\pm$ 0.002 0.851 $\pm$ 0.018 0.822 $\pm$ 0.002 0.805 $\pm$ 0.002 0.822 $\pm$ 0.002
20 Time 1.616 $\pm$ 0.036 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 2.057 $\pm$ 0.011 1.171 $\pm$ 0.008
RMSE 0.255 $\pm$ 0.006 0.254 $\pm$ 0.005 0.254 $\pm$ 0.005 0.554 $\pm$ 0.070 0.253 $\pm$ 0.005 0.261 $\pm$ 0.005 0.254 $\pm$ 0.005
NLPD 0.847 $\pm$ 0.002 0.784 $\pm$ 0.002 0.822 $\pm$ 0.001 0.986 $\pm$ 0.053 0.821 $\pm$ 0.001 0.799 $\pm$ 0.002 0.822 $\pm$ 0.001
Table 3: Results of the aggregation models for SVGP on the UCI datasets of size $\sim$45K. Best and worst performances are highlighted in yellow and red, respectively.
Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
Kin40k 8 Time 1.821 ± 0.192 5.077 ± 0.110 5.077 ± 0.110 5.077 ± 0.110 5.077 ± 0.110 7.513 ± 0.158 5.077 ± 0.110
RMSE 0.391 ± 0.030 1.006 ± 0.007 1.006 ± 0.007 1.510 ± 0.128 0.997 ± 0.007 0.471 ± 0.016 0.906 ± 0.007
NLPD 0.979 ± 0.014 1.440 ± 0.009 1.469 ± 0.005 1.873 ± 0.120 1.463 ± 0.005 0.971 ± 0.011 1.322 ± 0.004
16 Time 1.821 ± 0.192 2.559 ± 0.014 2.559 ± 0.014 2.559 ± 0.014 2.559 ± 0.014 4.425 ± 0.019 2.559 ± 0.014
RMSE 1.233 ± 0.300 1.008 ± 0.009 1.008 ± 0.009 2.639 ± 0.384 0.997 ± 0.007 0.483 ± 0.010 0.816 ± 0.008
NLPD 1.735 ± 0.269 1.455 ± 0.013 1.452 ± 0.006 3.591 ± 0.765 1.444 ± 0.004 0.929 ± 0.007 1.176 ± 0.005
64 Time 1.821 ± 0.192 1.062 ± 0.012 1.062 ± 0.012 1.062 ± 0.012 1.062 ± 0.012 1.305 ± 0.009 1.062 ± 0.012
RMSE 1.233 ± 0.300 0.997 ± 0.007 0.997 ± 0.007 1.902 ± 0.554 0.997 ± 0.007 0.611 ± 0.038 0.727 ± 0.005
NLPD 1.735 ± 0.269 1.448 ± 0.010 1.469 ± 0.005 2.351 ± 0.715 1.469 ± 0.005 1.011 ± 0.033 1.071 ± 0.003
KEGG 8 Time 1.818 ± 0.256 6.211 ± 0.068 6.211 ± 0.068 6.211 ± 0.068 6.211 ± 0.068 9.227 ± 0.085 6.211 ± 0.068
RMSE 1.152 ± 0.162 1.408 ± 0.034 1.408 ± 0.034 2.527 ± 0.910 1.378 ± 0.008 1.096 ± 0.218 1.267 ± 0.031
NLPD 1.572 ± 0.151 1.969 ± 0.061 1.769 ± 0.021 2.900 ± 1.259 1.750 ± 0.008 1.567 ± 0.280 1.592 ± 0.019
16 Time 1.818 ± 0.256 3.449 ± 0.041 3.449 ± 0.041 3.449 ± 0.041 3.449 ± 0.041 5.407 ± 0.039 3.449 ± 0.041
RMSE 1.401 ± 0.062 1.455 ± 0.043 1.455 ± 0.043 7.284 ± 2.053 1.378 ± 0.008 1.355 ± 0.383 1.178 ± 0.035
NLPD 1.841 ± 0.071 2.147 ± 0.087 1.797 ± 0.027 14.708 ± 7.131 1.745 ± 0.004 2.072 ± 0.735 1.455 ± 0.022
64 Time 1.818 ± 0.256 1.095 ± 0.011 1.095 ± 0.011 1.095 ± 0.011 1.095 ± 0.011 2.075 ± 0.015 1.095 ± 0.011
RMSE 1.401 ± 0.062 1.386 ± 0.010 1.386 ± 0.010 9.475 ± 2.704 1.378 ± 0.008 1.250 ± 0.210 1.011 ± 0.007
NLPD 1.841 ± 0.071 2.088 ± 0.021 1.762 ± 0.004 21.055 ± 11.376 1.757 ± 0.006 1.882 ± 0.346 1.285 ± 0.003
Protein 8 Time 1.876 ± 0.181 6.113 ± 0.103 6.113 ± 0.103 6.113 ± 0.103 6.113 ± 0.103 8.637 ± 0.136 6.113 ± 0.103
RMSE 0.773 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.783 ± 0.013 0.773 ± 0.003 0.773 ± 0.003 0.773 ± 0.003
NLPD 1.222 ± 0.002 1.176 ± 0.003 1.279 ± 0.002 1.285 ± 0.008 1.279 ± 0.002 1.206 ± 0.004 1.279 ± 0.002
16 Time 1.876 ± 0.181 3.484 ± 0.034 3.484 ± 0.034 3.484 ± 0.034 3.484 ± 0.034 5.427 ± 0.052 3.484 ± 0.034
RMSE 0.774 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.795 ± 0.022 0.773 ± 0.003 0.774 ± 0.003 0.773 ± 0.003
NLPD 1.222 ± 0.002 1.171 ± 0.003 1.279 ± 0.002 1.293 ± 0.013 1.279 ± 0.002 1.188 ± 0.005 1.279 ± 0.002
64 Time 1.876 ± 0.181 1.120 ± 0.024 1.120 ± 0.024 1.120 ± 0.024 1.120 ± 0.024 2.155 ± 0.060 1.120 ± 0.024
RMSE 0.774 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.893 ± 0.100 0.773 ± 0.003 0.776 ± 0.008 0.773 ± 0.003
NLPD 1.222 ± 0.002 1.168 ± 0.003 1.278 ± 0.002 1.360 ± 0.070 1.278 ± 0.002 1.175 ± 0.008 1.278 ± 0.002

### v2

@@ 5.1.2 Prediction
Figure 4 compares the performance of various aggregation models for exact GPs using four metrics: computational time, root mean squared error (RMSE), negative log predictive density (NLPD), and 2-Wasserstein distance to the full exact GP (measuring how closely each aggregated model approximates the full GP). While grBCM performs the best in RMSE, NLPD, and 2-Wasserstein distance, it requires more computational time due to the larger size of the augmented local training datasets and has larger standard deviations of the metrics. Our method, GP with optimal weights, provides comparable predictions with significantly lower computational cost. Figure 5 also compares four metrics for SVGP across all aggregation models. It’s observed that our method and grBCM significantly outperform other methods for SVGP.
5.2 UCI regression for temporal extrapolation
In this section, we evaluate the aggregation models on three real-world time-series datasets from the UCI repository (Asuncion and Newman, 2007): Gas Turbine CO and NOx Emission (Kaya et al., 2019), Beijing PM2.5 (Liang et al., 2015), and Metro Interstate Traffic Volume (Hogue, 2019). Unlike standard random train-test splits, we employ temporal extrapolation splits where models are trained on earlier years and tested on future years, creating a more realistic and challenging evaluation scenario that tests the models’ ability to handle distribution shift and regime changes.
Dataset descriptions and splits:
  • Gas Turbine CO & NOx (36,733 samples, 10 features): Hourly sensor measurements from 2011–2015. Training on 2011–2013, testing on 2014–2015. This dataset contains two prediction targets: CO and NOx emissions, we run experiments independently for each target. The future years contain operational regimes not fully represented in training data.
  • Beijing PM2.5 (43,824 samples, 13 features): Hourly air quality measurements from 2010–2014. Training on 2010–2013, testing on 2014. The test period includes extreme pollution events such as haze episodes that challenge model calibration.
  • Metro Traffic (48,204 samples, 8 features): Hourly traffic volume measurements from 2012–2018. Training on 2012–2016, testing on 2017–2018. Traffic patterns exhibit regime shifts from construction, policy changes, and seasonal variations.
Experimental setup:
To create a challenging distributed setting where expert diversity and correlation become important, we partition each training dataset into $M=20,40,60$ heterogeneous experts by randomly shuffling and dividing the data. Specifically, we use non-shared hyperparameters across experts (i.e., each expert optimizes its own lengthscale, kernel variance, and noise variance independently), which creates the strong correlated local datasets that our method can handle. We employ SVGP with $m=128$ inducing points and RBF kernel. All aggregation models utilize FedAvg with the Adam optimizer, with 500 training iterations to ensure sufficient convergence for the extrapolation task. Results are averaged over 10 different random seeds.
Figure 6 presents the performance metrics of the aggregation models on these four datasets. It’s observed that: (1) BCM and rBCM perform poorly, particularly for NOx and Beijing PM2.5, where it yields significantly degraded RMSE and NLPD due to overconfident but incorrect experts being equally weighted. (2) While grBCM shows competitive RMSE, it requires substantially more computation time due to the need to train additional models. (3) Our method, distributed SVGP with optimal weights (opt), consistently achieves the best or near-best RMSE and NLPD across all datasets while maintaining computational efficiency comparable to simpler baselines. Notably, on the challenging extrapolation tasks (NOx and Metro Traffic), our method’s ability to account for expert correlation and reliability provides clear advantages in both prediction accuracy and uncertainty quantification. Detailed metric values are reported in Table 2.
Figure 6: Comparison of aggregation models for SVGP on temporal extrapolation tasks. Models are trained on earlier years and tested on future years to evaluate performance under distribution shift.
Table 2: Results of aggregation models for SVGP on UCI datasets for temporal extrapolation tasks. Models are trained on earlier years and tested on future years. Best and worst performances are highlighted in yellow and red, respectively.
Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
Gas Turbine CO 20 Time 35.405 $\pm$ 0.415 0.164 $\pm$ 0.002 0.164 $\pm$ 0.001 0.163 $\pm$ 0.001 0.164 $\pm$ 0.001 0.297 $\pm$ 0.001 0.163 $\pm$ 0.001
RMSE 1.642 $\pm$ 0.053 1.518 $\pm$ 0.005 1.518 $\pm$ 0.005 2.828 $\pm$ 0.054 2.510 $\pm$ 0.045 1.615 $\pm$ 0.117 1.519 $\pm$ 0.005
NLPD 1.955 $\pm$ 0.047 2.310 $\pm$ 0.009 2.157 $\pm$ 0.007 5.881 $\pm$ 0.195 4.350 $\pm$ 0.123 1.991 $\pm$ 0.199 1.830 $\pm$ 0.004
40 Time 35.405 $\pm$ 0.415 0.100 $\pm$ 0.008 0.095 $\pm$ 0.005 0.094 $\pm$ 0.005 0.093 $\pm$ 0.005 0.157 $\pm$ 0.005 0.091 $\pm$ 0.005
RMSE 1.656 $\pm$ 0.000 1.526 $\pm$ 0.002 1.525 $\pm$ 0.002 3.106 $\pm$ 0.020 2.939 $\pm$ 0.019 1.733 $\pm$ 0.214 1.527 $\pm$ 0.004
NLPD 1.973 $\pm$ 0.000 2.366 $\pm$ 0.005 2.273 $\pm$ 0.006 7.269 $\pm$ 0.085 6.170 $\pm$ 0.071 2.244 $\pm$ 0.378 1.833 $\pm$ 0.004
60 Time 35.405 $\pm$ 0.415 0.079 $\pm$ 0.007 0.074 $\pm$ 0.005 0.072 $\pm$ 0.002 0.072 $\pm$ 0.002 0.109 $\pm$ 0.001 0.069 $\pm$ 0.001
RMSE 1.656 $\pm$ 0.000 1.550 $\pm$ 0.004 1.549 $\pm$ 0.005 3.187 $\pm$ 0.021 3.078 $\pm$ 0.020 1.789 $\pm$ 0.196 1.552 $\pm$ 0.005
NLPD 1.973 $\pm$ 0.000 2.432 $\pm$ 0.010 2.360 $\pm$ 0.010 7.748 $\pm$ 0.094 6.935 $\pm$ 0.077 2.370 $\pm$ 0.261 1.849 $\pm$ 0.005
Gas Turbine NOx 20 Time 36.239 $\pm$ 1.322 0.164 $\pm$ 0.003 0.165 $\pm$ 0.003 0.164 $\pm$ 0.003 0.165 $\pm$ 0.003 0.299 $\pm$ 0.008 0.164 $\pm$ 0.003
RMSE 1.065 $\pm$ 0.041 1.070 $\pm$ 0.012 1.070 $\pm$ 0.012 1.964 $\pm$ 0.035 1.949 $\pm$ 0.032 1.095 $\pm$ 0.083 1.066 $\pm$ 0.012
NLPD 4.103 $\pm$ 0.446 7.900 $\pm$ 0.171 7.898 $\pm$ 0.176 26.145 $\pm$ 0.892 25.703 $\pm$ 0.802 5.297 $\pm$ 1.285 4.032 $\pm$ 0.100
40 Time 36.239 $\pm$ 1.322 0.104 $\pm$ 0.007 0.099 $\pm$ 0.006 0.095 $\pm$ 0.004 0.094 $\pm$ 0.004 0.162 $\pm$ 0.008 0.092 $\pm$ 0.004
RMSE 1.115 $\pm$ 0.000 1.063 $\pm$ 0.008 1.063 $\pm$ 0.008 2.036 $\pm$ 0.017 2.021 $\pm$ 0.016 1.096 $\pm$ 0.057 1.062 $\pm$ 0.013
NLPD 4.332 $\pm$ 0.000 7.990 $\pm$ 0.126 7.987 $\pm$ 0.122 28.948 $\pm$ 0.476 28.460 $\pm$ 0.443 5.350 $\pm$ 0.713 3.890 $\pm$ 0.101
60 Time 36.240 $\pm$ 1.322 0.084 $\pm$ 0.003 0.076 $\pm$ 0.005 0.075 $\pm$ 0.004 0.075 $\pm$ 0.003 0.112 $\pm$ 0.003 0.072 $\pm$ 0.002
RMSE 1.115 $\pm$ 0.000 1.066 $\pm$ 0.007 1.066 $\pm$ 0.007 2.085 $\pm$ 0.020 2.069 $\pm$ 0.020 1.043 $\pm$ 0.102 1.065 $\pm$ 0.007
NLPD 4.332 $\pm$ 0.000 8.104 $\pm$ 0.112 8.100 $\pm$ 0.111 30.656 $\pm$ 0.594 30.121 $\pm$ 0.565 5.759 $\pm$ 1.683 3.860 $\pm$ 0.053
Beijing PM2.5 20 Time 60.044 $\pm$ 1.202 0.242 $\pm$ 0.007 0.240 $\pm$ 0.007 0.240 $\pm$ 0.007 0.240 $\pm$ 0.007 0.422 $\pm$ 0.008 0.239 $\pm$ 0.007
RMSE 7.780 $\pm$ 0.116 7.776 $\pm$ 0.016 7.775 $\pm$ 0.015 8.530 $\pm$ 0.048 8.363 $\pm$ 0.039 7.755 $\pm$ 0.094 7.775 $\pm$ 0.018
NLPD 8.640 $\pm$ 0.652 35.010 $\pm$ 0.188 30.668 $\pm$ 0.299 38.362 $\pm$ 0.289 31.270 $\pm$ 0.280 9.147 $\pm$ 0.602 7.350 $\pm$ 0.100
40 Time 60.044 $\pm$ 1.202 0.138 $\pm$ 0.008 0.127 $\pm$ 0.003 0.125 $\pm$ 0.002 0.125 $\pm$ 0.002 0.228 $\pm$ 0.002 0.124 $\pm$ 0.002
RMSE 7.655 $\pm$ 0.000 7.830 $\pm$ 0.013 7.829 $\pm$ 0.013 8.823 $\pm$ 0.055 8.682 $\pm$ 0.054 7.786 $\pm$ 0.119 7.831 $\pm$ 0.010
NLPD 9.020 $\pm$ 0.000 39.237 $\pm$ 0.137 36.061 $\pm$ 0.179 46.846 $\pm$ 0.554 40.320 $\pm$ 0.481 11.787 $\pm$ 2.199 7.232 $\pm$ 0.096
60 Time 60.044 $\pm$ 1.202 0.123 $\pm$ 0.016 0.095 $\pm$ 0.003 0.094 $\pm$ 0.002 0.095 $\pm$ 0.003 0.159 $\pm$ 0.004 0.094 $\pm$ 0.002
RMSE 7.655 $\pm$ 0.000 7.871 $\pm$ 0.012 7.869 $\pm$ 0.012 9.041 $\pm$ 0.055 8.917 $\pm$ 0.047 7.942 $\pm$ 0.157 7.873 $\pm$ 0.013
NLPD 9.020 $\pm$ 0.000 41.091 $\pm$ 0.128 38.502 $\pm$ 0.151 51.722 $\pm$ 0.562 45.666 $\pm$ 0.351 12.596 $\pm$ 1.516 7.140 $\pm$ 0.095
Metro Traffic 20 Time 57.739 $\pm$ 2.806 0.317 $\pm$ 0.051 0.262 $\pm$ 0.010 0.255 $\pm$ 0.010 0.254 $\pm$ 0.009 0.467 $\pm$ 0.018 0.253 $\pm$ 0.010
RMSE 6.065 $\pm$ 0.701 5.612 $\pm$ 0.099 5.534 $\pm$ 0.060 58.979 $\pm$ 2.695 34.593 $\pm$ 0.416 6.512 $\pm$ 1.272 5.741 $\pm$ 0.199
NLPD 7.874 $\pm$ 0.083 11.481 $\pm$ 0.157 7.960 $\pm$ 0.023 235.124 $\pm$ 9.921 30.309 $\pm$ 1.069 7.942 $\pm$ 0.230 7.841 $\pm$ 0.026
40 Time 57.739 $\pm$ 2.806 0.419 $\pm$ 0.118 0.162 $\pm$ 0.002 0.151 $\pm$ 0.004 0.146 $\pm$ 0.001 0.236 $\pm$ 0.003 0.144 $\pm$ 0.001
RMSE 6.236 $\pm$ 0.000 5.587 $\pm$ 0.027 5.569 $\pm$ 0.020 49.724 $\pm$ 0.474 41.100 $\pm$ 0.278 6.050 $\pm$ 0.663 5.648 $\pm$ 0.067
NLPD 7.893 $\pm$ 0.001 16.096 $\pm$ 0.074 9.088 $\pm$ 0.020 354.159 $\pm$ 3.765 76.711 $\pm$ 0.896 7.872 $\pm$ 0.145 7.843 $\pm$ 0.012
60 Time 57.740 $\pm$ 2.806 0.330 $\pm$ 0.128 0.137 $\pm$ 0.004 0.122 $\pm$ 0.002 0.122 $\pm$ 0.002 0.167 $\pm$ 0.006 0.118 $\pm$ 0.001
RMSE 6.236 $\pm$ 0.000 5.651 $\pm$ 0.023 5.638 $\pm$ 0.020 44.833 $\pm$ 0.618 40.345 $\pm$ 0.422 5.884 $\pm$ 0.406 5.692 $\pm$ 0.069
NLPD 7.893 $\pm$ 0.001 20.677 $\pm$ 0.092 10.498 $\pm$ 0.038 433.313 $\pm$ 6.898 113.021 $\pm$ 0.620 7.829 $\pm$ 0.143 7.856 $\pm$ 0.012
6 Discussions
Conclusions
@@
In this work, we introduced an algorithm to enhance GP regressions with inducing points on sparse grids using the optimized combination technique (OptiCom). Building on this foundation, we proposed a novel aggregated prediction algorithm suitable for both distributed exact GP and SVGP. This method leverages predictions from local models and optimally combines them using weights derived from solving a linear system. Additionally, it inherently incorporates correlations between experts through the computation of optimal weights. The proposed methods have a computational complexity requiring $\mathcal{O}(M^{3})$ additional time compared to PoE family and BCM family for $M$ GP experts. However, this overhead is minimal in practice, as the number of experts $M$ is typically small compared to the size of the local datasets. Through extensive experiments and comparisons with state-of-the-art aggregation models, we demonstrated the competitiveness and stability of our proposed method. Our algorithm consistently provides stable and accurate predictions, offering a robust solution for distributed GP regression.
Practical considerations
Beyond model design, practical deployment involves key trade-offs in choosing the number of experts and inducing points. The choice of the number of experts $M$ involves a trade-off between efficiency and accuracy. Increasing $M$ improves computational efficiency, as each expert handles a smaller subset of the data and can be trained in parallel. However, if $M$ becomes too large, each expert receives too little data, degrading the quality of the local GP approximations. This effect is clearly illustrated in Figure 4. In practice, we recommend selecting $M$ based on empirical validation, such as cross-validation. A practical set of initial candidates can be defined as $M\in\{\left\lfloor\sqrt{n}/i\right\rfloor:i=4,5,6\}$, where $n$ is size of the global training data.
Similarly, the number of inducing points affects both accuracy and computational cost. More inducing points improve local approximations, and consequently lead to more accurate aggregated predictions, they also increase computational cost. In practice, selecting 5–10% of the local training data size as the number of inducing points provides a good balance.
Limitations
While our method guarantees consistent predictive means in the limit, it does not ensure consistent predictive variances. Theoretically, variance consistency would require the aggregation weights to satisfy an additional $\ell_{2}$-norm condition, which we do not impose. As a result, variance estimates may be slightly less accurate compared to methods like grBCM, especially in uncertainty-sensitive applications.
Future directions
A direction for future work is to extend our framework such that it also ensures variance consistency, potentially by modifying the weight optimization process or incorporating additional constraints. Furthermore, integrating adaptive selection of expert sizes and inducing points could further enhance the scalability and flexibility of the method in large-scale settings.
Acknowledgements
@@
Data availability statement
The data that support the findings of this study are openly available in (Asuncion and Newman, 2007) at https://archive.ics.uci.edu/datasets.
@@ Appendix A GP with OptiCom
In this section, we provide supporting figures for optimized combination technique (OptiCom) and detailed algorithms for GP with OptiCom in Section A.1 and Section A.2 respectively.
@@ A.2 Algorithms
4:for $t=0$ to b1b-1 do # for loop over $t$ and $i$ can be parallelized
5:  for $i=1$ to btb-t do # loop over the diagonal with offset =t=t
@@
[…]
@@
[…]
@@
[…]
@@ Appendix B Distributed GP with optimal weights
In this section, we outline the algorithms of the proposed aggregation model for distributed SVGP and distributed exact GP in Section B.1 and Section B.2 respectively.
B.1 SVGP
@@
[…]
6:for $t=0$ to M1M-1 do # for loop over $t$ and $i$ can be parallelized
7:  for $i=1$ to MtM-t do # loop over the diagonal with offset =t=t
@@
[…]
@@
[…]
@@ B.2 Exact GP
5:for $t=0$ to M1M-1 do # for loop over $t$ and $i$ can be parallelized
6:  for $i=1$ to MtM-t do # loop over the diagonal with offset =t=t
@@
[…]
@@
[…]
@@ Appendix C Experiment
In this section, we provide the supporting boxplots for the distributed training on synthetic data and tables for the aggregation models on UCI datasets in Section C.1.
@@ C.1 Boxplots of the hyperparameter estimates
Figure 8: Boxplots of the optimized hyperparameter estimates on the $n=10^{4}$ dataset.